OpenStudy (anonymous):

solve -16less than or equal3x-9less than or equal-2

OpenStudy (anonymous):

-16<=3x-9<=-2

OpenStudy (anonymous):

-16 <= 3x-9 =< -2 add 9 everywhere; -7 <= 3x <= 7 divide all by 3; -7/3 <= x <= 7/3 think thats how its done anyways

OpenStudy (anonymous):

-7/3<=x<=7/3